OpenStudy (fanduekisses):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (fanduekisses):

d=42 right

hartnn (hartnn):

yes, thats correct

hartnn (hartnn):

find a1 = first term now use, \(\Large a_n = a_1 +(n-1)d\) plug in n = 15, a15 = 9 , d= 42

OpenStudy (fanduekisses):

so 9=a1 + (15-1)42 solve a1= -579 so the equation is an=-579 + (n-1)42 ^_^

hartnn (hartnn):

yes! thats corect :) if you want you can simplify that :)

OpenStudy (fanduekisses):

thanks :)